OpenStudy (anonymous):

7.Salisbury High School decided to take their students on a field trip to a theme park. A total of 150 people went on the trip. Adults pay $45.00 for a ticket and students pay $28.50 for a ticket. How many students and how many adults went to the park if they paid a total of $4770?

OpenStudy (anonymous):

a+s=150 a*45+28.5s=4770 :D that is your system

OpenStudy (anonymous):

so how do you solve that equation to find out the answer? that's the problem im having. I know how to set it up but how to solve is the problem.

OpenStudy (anonymous):

that is easy part lol. a+s=150 a*45+28.5s=4770 well lets solve for a in eq 1 a=150-s we plug it into eq2 (150-s)*45+28.5s=4770 now we solve for s (150-s)*45+28.5*s=4770 6750-45s+28.5*s= 4700 expand left side particularly this (150-s)*45 6750-16.5s=4700 now we add like terms the -45s+28.5*s -16.5s =-2050 now we subtract 7550 from both sides s=124.24 students looks odd tho se we will say 124 students

OpenStudy (anonymous):

a+124=150 a=150-124 = 26 adults